OpenStudy (anonymous):

how would i graph this? |x + 5| y – 2

OpenStudy (anonymous):

try wolframalpha.com

OpenStudy (anonymous):

yeah.. thats not really helping its an absolute value inequality im sorry

OpenStudy (anonymous):

\[\left| x+5 \right| \ge y-2\]

OpenStudy (anonymous):

@whpalmer4

OpenStudy (whpalmer4):

Well, you can break this down into two sections: the section where (x+5) is positive, and the section where (x+5) is negative. Can you graph \[x+5\ge y-2\]?

OpenStudy (whpalmer4):

That simplifies to \[x\ge y-7\]or \[-y \ge -x-7\]\[y \le x+7\]